OpenStudy (johnoneill):

(6/z^2)-(8/z) what do i do help

OpenStudy (sshayer):

\[\frac{ 6 }{ z^2 }-\frac{ 8 }{ z }=\frac{ 6-8z }{ z^2 }\]